OpenStudy (anonymous):

Solve for p and q: -3+2p=5-7q -1+p=-2q -25-8p=3-2q Following multiplication of one equation, is that equation being subtracted or subtracts the other?

OpenStudy (anonymous):

-3+2p=5-7q--(1) -1+p=-2q--(2) -25-8p=3-2q--(3) (3)-(2)=>-24-9p=3 9p=-27 p=-3 -2q=-1-3 -2q=-4 q=2

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw